# CoordField: Coordination Field for Agentic UAV Task Allocation In Low-altitude Urban Scenarios

### v1

@@ II-A Agentic Systems
In recent years, Agentic AI has emerged as a novel intelligent paradigm characterized by autonomous perception, goal-driven planning, and adaptive decision-making, offering innovative and effective solutions for multi-agent collaborative tasks [8, 9]. Unlike traditional passive approaches, Agentic AI emphasizes continuous goal pursuit, contextual memory, adaptive tool use, and self-reflection capabilities of agents [10]. These abilities enable agents to operate with greater autonomy and coordination in dynamic environments. Agent systems powered by LLMs have demonstrated strong reasoning and planning capabilities, particularly in complex task decomposition and sequential execution [3]. These features provide a solid theoretical and methodological foundation for building autonomous multi-UAV systems. In complex urban task scenarios [11], Agentic AI can assist UAVs in understanding high-level instructions, distributing subtasks effectively, and rapidly reconfiguring coordination strategies in response to task changes or agent failures, thereby enabling truly autonomous collaboration.
@@ IV-B Experimental Results and Analysis
The experimental evaluation on 50 natural language command tasks demonstrates that the proposed coordination field agentic system outperforms all baseline models across multiple key performance metrics as shown in Table 1. Specifically, it achieves the highest TPA of 96%, significantly exceeding GPT-4o (74%)[20], Claude-3-7-Sonnet (76%)[21], and Deepseek-v3 (70%)[22], indicating its superior semantic understanding and robustness in converting unstructured commands into executable UAV instructions. In terms of CE, the system maintains an average rate of 95%, consistently outperforming other models such as GPT-4o (85%) and LLaMA-4-Scout (79%)[23], and remains highly reliable even under task-burst scenarios with over 80% effective coverage. With respect to TLB, our method demonstrates the lowest task distribution deviation of only 0.8, compared to 2.1 and 1.6 observed in Deepseek-v3 and Claude-3-7-Sonnet respectively, which significantly reduces resource clustering and enhances spatial coordination. Additionally, the UUR reaches 97%, confirming that the majority of UAVs remain actively engaged in task execution, whereas competing models such as GPT-4o and LLaMA-4-Scout fall below 85%. These results collectively validate the system’s strong coordination capability, dynamic adaptability, and overall efficiency in managing heterogeneous UAV swarms within complex urban environments.
@@
This paper proposes a coordination field Agentic system for UAV task coordination. The system demonstrates efficiency, stability, and adaptability, enabling real-time parsing of natural language instructions and efficient task allocation via potential fields, fluid-based velocity fields, and vortex-based guidance. Experimental results verify its superior performance in complex and dynamic urban environments. In future work, we plan to extend the system to three-dimensional scenarios and real-world UAV swarms.

### v2

@@ II-A Agentic Systems
In recent years, Agentic AI has emerged as a novel intelligent paradigm characterized by autonomous perception, goal-driven planning, and adaptive decision-making, offering innovative and effective solutions for multi-agent collaborative tasks [8, 9]. Unlike traditional passive approaches, Agentic AI emphasizes continuous goal pursuit, contextual memory, adaptive tool use, and self-reflection capabilities of agents [10]. These abilities enable agents to operate with greater autonomy and coordination in dynamic environments. Agent systems powered by LLMs have demonstrated strong reasoning and planning capabilities, particularly in complex task decomposition and sequential execution [3]. These features provide a solid theoretical and methodological foundation for building autonomous multi-UAV systems. In complex urban task scenarios [11], Agentic AI can assist UAVs in understanding high-level instructions, distributing subtasks effectively, and rapidly reconfiguring coordination strategies in response to task changes or agent failures, thereby enabling truly autonomous collaboration [12, 13].
@@ IV-B Experimental Results and Analysis
The experimental evaluation on 50 natural language command tasks demonstrates that the proposed coordination field agentic system outperforms all baseline models across multiple key performance metrics as shown in Table 1. Specifically, it achieves the highest TPA of 96%, significantly exceeding GPT-4o (74%), Claude-3-7-Sonnet (76%), and Deepseek-v3 (70%), indicating its superior semantic understanding and robustness in converting unstructured commands into executable UAV instructions. In terms of CE, the system maintains an average rate of 95%, consistently outperforming other models such as GPT-4o (85%) and LLaMA-4-Scout (79%), and remains highly reliable even under task-burst scenarios with over 80% effective coverage. With respect to TLB, our method demonstrates the lowest task distribution deviation of only 0.8, compared to 2.1 and 1.6 observed in Deepseek-v3 and Claude-3-7-Sonnet respectively, which significantly reduces resource clustering and enhances spatial coordination. Additionally, the UUR reaches 97%, confirming that the majority of UAVs remain actively engaged in task execution, whereas competing models such as GPT-4o and LLaMA-4-Scout fall below 85%. These results collectively validate the system’s strong coordination capability, dynamic adaptability, and overall efficiency in managing heterogeneous UAV swarms within complex urban environments.
@@
This paper proposes a coordination field Agentic system for UAV task coordination. The system demonstrates high efficiency, stability and adaptability, and is able to parse natural language instructions in real time and achieve efficient task allocation through coordination field guidance. Experimental results verify its superior performance in complex and dynamic urban environments. In future work, we plan to extend the system to three-dimensional scenarios and real-world UAV swarms[22, 23].